### v1

# Your Classifier Can Do More: Towards Bridging the Gaps in Classification, Robustness, and Generation

Abstract
Joint Energy-based Models (JEMs), a class of hybrid generative-discriminative models, are well known for their ability to achieve both high classification accuracy and generative capability within a single model. However, their robustness still lags significantly behind the classifiers based adversarial training (AT). Conversely, while AT is currently the most effective approach to improving the classifier’s robustness, it typically sacrifices accuracy on clean data and lacks generative capability. The triple trade-off between classification accuracy, generative capability and robustness, raises a natural question: Can a single model simultaneously achieve high classification accuracy, adversarial robustness, and generative performance? – a goal that has been rarely explored. To address this question, we systematically analyze the energy distribution differences of clean, adversarial, and generated samples across various JEM variants and adversarially trained models. We observe that AT tends to reduce the energy gap between clean and adversarial samples, while JEMs reduce the gap between clean and synthetic ones. This observation suggests a key insight: if the energy distributions of all three data types can be aligned, we might unify the strengths of AT and JEMs, resolving their inherent trade-offs. Building on this idea, we propose Energy-based Joint Distribution Adversarial Training (EB-JDAT), to jointly model the clean data distribution, the adversarial distribution, and the classifier by maximizing their joint probability. To capture the full adversarial distribution, we introduce a new min-max energy optimization strategy. EB-JDAT is a general and flexible optimization method, compatible with various JEM variants. Extensive experimental results demonstrate that EB-JDAT not only maintains near original accuracy and generative capability of JEMs, but also significantly enhances robustness, even surpassing state-of-the-art ATs.
1 Introduction
Recently, Joint Energy-based Models (JEMs) grathwohl2019your ; yang2021jem++ ; yang2023towards have garnered significant attention for their ability to integrate generative modeling within a discriminative framework. This approach effectively bridges classification and generation from an energy-based perspective and exhibits surprising inherent robustness. However, JEMs still exhibit limited robustness compared to Adversarial Training (AT) mkadry2017towards . On the other hand, AT jia2022adversarial ; hammoudeh2024provable ; bortolussi2024robustness ; fang20253sat is widely recognized as the most effective method for improving the robustness of discriminative classifiers. However, it typically sacrifices accuracy on clean data and lacks generative capability zhang2019theoretically . Consequently, the inherent gap in classification, robustness and generation raises a natural question: Can a single model simultaneously achieve high classification accuracy, adversarial robustness, and generative performance? – a goal that has been rarely explored.
Figure 1: Distributions of the $E_{\theta}(\textbf{x},y)$ of adversarial samples for PGD-20 vs. generated samples vs. clean samples on CIFAR-10. indicates generated samples, indicates adversarial samples, indicates clean samples.
Table 1: Mean and variance of one-to-one energy differences between clean and adv.
[…]
Standard 10.18 3.27 0.06
PGD-AT 1.46 0.83 55.08
TRADES 1.01 0.45 56.10
JEM 5.93 7.30 7.28
SADAJEM 2.66 3.83 42.46
EB-JDAT 0.89 0.31 66.12
To answer this question, we begin by investigating, from an energy perspective, the underlying causes of the performance gap between AT mkadry2017towards ; zhang2019theoretically and JEMs grathwohl2019your ; yang2023towards in terms of accuracy, generative capability, and adversarial robustness. We compute the energy distributions of clean vs. adversarial vs. generated samples by utilizing different models, and the results are shown in Fig. 1. First, the energy distributions of clean vs. adversarial samples, give a clear indication of AT tends to reduce the energy gap between clean and adversarial samples. Interestingly, JEMs implicitly reduce the gap between them, even though the distribution of adversarial samples is not explicitly incorporated during training, which explains the potential source of robustness in JEMs. Further, we compute the mean and variance of one-to-one energy differences between them. As reported in Tab. 1, models with smaller energy differences exhibit higher robustness. Next, we investigate the classification and generation gaps by analyzing the energy distributions of clean vs. generated samples in Fig. 1. Compared to AT, JEMs reduce the gap between them, achieving high accuracy in classification while maintaining strong generative quality. Building on these observations, we draw a key insight: if the energy distributions of all three data types can be aligned, we might unify the strengths of AT and JEMs, resolving their inherent trilemma.
Inspired by this idea, we align the energy distribution of the three data types by jointly modeling the clean data distribution, the adversarial distribution and the classifier, by learning $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$ via energy-based model, where $\mathbf{x}$, $y$ and $\tilde{\mathbf{x}}$ denote the clean samples, labels and adversarial examples, and $\theta$ denote the model parameters. Through Bayesian decomposition, $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$ can be further factorized as the data distribution $p_{\theta}(\mathbf{x})$, adversarial distribution $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$ and $p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$. While $p_{\theta}(\mathbf{x})$ can be computed via approximated sampling methods nijkamp2020anatomy and $p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$ is simply a cross-entropy objective for robust classification, computing $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$ is not straightforward because the full adversarial distribution is not observed during the training phase. However, our key observation is that adversarial perturbations almost surely move samples out of the original data manifold, which lies in high-density areas, into low-density areas where the classifier is more likely to fail. Motivated by this observation, we propose a new min-max energy optimization to approximate $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$. During the sampling phase, we search for high-energy adversarial examples, then minimize the energy gap between adversarial and clean samples during the training phase, thereby pulling the adversarial samples back into low-energy regions. We name our proposed method as Energy-based Joint Distribution Adversarial Training, or EB-JDAT. Extensive experiments across multiple datasets and attacks show that EB-JDAT achieves 64.76% and 35.63% robustness (under AutoAttack) on CIFAR-10 and CIFAR-100 datasets, outperforming the SOTA AT methods by +9.24% and +4.86% respectively, while maintaining near original clean accuracy and generative capability, which is competitive to most advanced JEMs.
2 Related work
Adversarial Defense.
Since the discovery of adversarial vulnerabilities in deep learning models yuan2021automa ; ma2023transferable , numerous defense strategies have been proposed to mitigate this security challenge. Prominent approaches include input denoising li2022feature , gradient regularization li2023revisiting ; jia2022prior , defensive distillation huang2023boosting , and adversarial training (AT) mkadry2017towards ; fang20253sat . Recent studies confirm that AT remains among the most effective defense methods sengphanith2023evaluating , significantly enhancing model robustness by incorporating adversarial examples into the training process. Notably, mkadry2017towards first formulated AT as a saddle-point optimization problem (PGD), laying the foundational framework for subsequent robust training methodologies. As adversarial attack methods have evolved, various novel strategies have emerged, refining conventional AT methods jia2022adversarial ; hammoudeh2024provable ; bortolussi2024robustness ; fang20253sat . However, many of these strategies prioritize robustness at the expense of clean accuracy. which is the principal evaluation metric in visual classification, highlighting an inherent trade-off in AT. To mitigate this issue, zhang2019theoretically proposed TRADES, which constrains the divergence in output probability between clean and adversarial examples. Additionally, co2022jacobian combined Jacobian regularization with model ensembling, introducing the Jacobians ensemble approach that effectively balances robustness and accuracy against universal adversarial perturbations. Furthermore, DUCAT wang2024new assigns a dummy class to each original class to absorb hard adversarial examples during training and remaps them back at inference to mitigating the trade-off. Although these approaches represent notable advancements, they still suffer from substantial accuracy degradation relative to standard classifiers, while offering only modest enhancements in robustness. Therefore, achieving a trade-off between accuracy and robustness remains an unresolved challenge in AT.
Robust Energy-based Model.
Recent studies have indicated that features extracted by robust classifiers are more closely aligned with human perceptual mechanisms compared to those derived from standard classifiers wangunified . The Joint Energy-based Model (JEM) grathwohl2019your reformulates conventional softmax classifiers into an energy-based model (EBM), facilitating hybrid discriminative-generative modeling and clarifying the connection between robustness and generative modeling. More recent efforts have focused on improving training efficiency and speed in JEM yang2021jem++ ; yang2023towards , effectively bridging the gap between classification and generation. Additionally, several approaches specifically focus on enhancing robustness, from an energy perspective. JEAT zhu2021towards initially demonstrated a link between AT and EBM, highlighting that despite differences in how these methods modify the energy function, they ultimately employ comparable contrastive strategies. Subsequently, mirza2024shedding reinterpreted robust classifiers within an energy-based framework and analyzed the energy landscape to reveal differential effects of targeted and non-targeted attacks during AT. However, these methods often sacrifice robustness in exchange for notable reductions in accuracy and generative capacity, with some even proving less robust than traditional adversarial defense strategies. Thus, bridging the inherent gap in classification, generation, and robustness is challenging, yet no existing method has effectively addressed this challenge. Although SOTA JEM variants perform well in classification and generative capabilities, their robustness remains limited.
3 Methodology
3.1 Preliminaries
Energy-based Models (EBMs)
From an energy perspective, the clean data distribution ${p}_{\theta}(\mathbf{x})$ can be explicitly parameterized through an energy function grathwohl2019your :
$p_{\theta}(\mathbf{x})=\frac{\exp\left(-E_{\theta}(\mathbf{x})\right)}{Z(\theta)}=\frac{\sum_{y\in\mathbf{y}}\exp\left(f_{\theta}(\mathbf{x})[y]\right)}{Z(\theta)},$ (1)
where $E_{\theta}(\cdot):\mathbb{R}^{D}\rightarrow\mathbb{R}$ denotes the energy function. This energy function assigns each data point a scalar value representing its density within the continuous data distribution. The normalizing constant $Z(\theta)=\int_{\mathbf{x}}\exp(-E_{\theta}(\mathbf{x}))$ ensures the proper normalization of probabilities. During training, the energy function $E_{\theta}(\mathbf{x})$ is optimized to assign lower energy values to high-density regions (e.g., training data) and higher energy values to low-density regions (e.g., adversarial samples). To optimize the energy-based model, a common strategy is to perform maximum likelihood estimation of the parameters $\theta$, in which the gradient of the log-likelihood is expressed as follows:
$\frac{\partial{log}p_{\theta}(\mathbf{x})}{\partial\theta}=\mathbb{E}_{p_{\theta}(\mathbf{x^{\prime}})}\left[\frac{\partial E_{\theta}(\mathbf{x^{\prime}})}{\partial\theta}\right]-\frac{\partial E_{\theta}(\mathbf{x})}{\partial\theta},$ (2)
where $\mathbf{x^{\prime}}$ is simply a sample obtained by sampling from the model distribution.
Joint Energy-based Models (JEMs)
JEM grathwohl2019your reinterprets the standard softmax classifier within the framework of an EBM. Specifically, it redefines the logits of the classifier $f_{\theta}$ to model the joint distribution $p_{\theta}(\textbf{x},y)$ of the input data x and their labels $y$ using an energy-based approach:
$p_{\theta}(\mathbf{x},y)=\frac{\exp\left(f_{\theta}(\mathbf{x})[y]\right)}{Z(\theta)}.$ (3)
In JEMs grathwohl2019your ; yang2021jem++ ; yang2023towards , the model parameters $\theta$ are optimized by maximizing $\log p_{\theta}(\mathbf{x},y)$:
$\log p_{{\theta}}(\mathbf{x},y)=\log p_{{\theta}}(y\mid\mathbf{x})+\log p_{{\theta}}(\mathbf{x}).$ (4)
Adversarial Training (AT)
Adversarial training can be viewed as a min-max optimization problem that integrates inner loss maximization and outer loss minimization mkadry2017towards . Given data $\mathbf{x}\in\mathbf{X}$, its corresponding label $y\in\mathbf{y}$, and a classification model $f_{\theta}$, the objective of the inner maximization is to identify the most aggressive adversarial examples by maximizing the loss. Subsequently, the outer minimization seeks to minimize the average loss of these adversarial examples. The specific optimization process is detailed in:
$\min_{\theta}\mathbb{E}_{(\mathbf{x},{y})\sim\mathcal{D}}\left[\max_{\|\tilde{\mathbf{x}}-\mathbf{x}\|\in\Omega}\mathbf{L}(\tilde{\mathbf{x}},{y};\theta)\right],$ (5)
where $\mathcal{D}$ denotes the training data, $\mathbf{L}$ represents the loss function (typically cross-entropy for classification), $\tilde{\mathbf{x}}$ is the adversarial example corresponding to $\mathbf{x}$, and $\Omega$ is the perturbation space.
3.2 Motivation
ATmkadry2017towards has demonstrated strong robustness against adversarial attacks. However, the adversarial samples introduced during training lie outside the true data manifold, causing the model to overemphasize these adversarial features while neglecting meaningful patterns in clean data ilyas2019adversarial . Consequently, AT-trained models exhibit a noticeable accuracy gap compared to standard classifiers. In contrast, JEMs have shown the ability to reinterpret discriminative classifiers as generative models grathwohl2019your , achieving notable performance in both image classification and generation. While JEMs exhibit improved adversarial robustness relative to standard models, a significant robustness gap still remains when compared to AT-trained methods. This naturally raises a new research question: Can we combine the strengths of AT and JEMs to develop a unified framework that mitigates, or even eliminates, triple trade-off between accuracy, robustness and generative performance?
To answer this question, we begin by visualizing the energy distributions of clean vs. adversarial vs. generated samples to investigate the source of the performance gains in JEMs and AT, which are shown in Fig. 1. We first analyze the energy distributions of clean vs. adversarial samples. Obviously, for models trained by AT, the energy distributions of clean and adversarial samples almost completely overlap. While for JEMs, the two distributions partially overlap but less than that observed in AT-based methods. Further, We compute the mean and variance of one-to-one energy differences between clean samples and their corresponding adversarial examples and report the results in Tab. 1. By analyzing these results, we have two findings: 1) the greater the overlap between the energy distributions of clean and adversarial samples, the higher the model’s robustness. 2) JEMs implicitly reduce the gap between adversarial and clean data, although the distribution of adversarial samples is not explicitly incorporated during training. This also explains the source of robustness in JEMs.
Next, we investigate the classification and generative gaps by analyzing the energy distributions of real vs. generated samples in Fig. 1. Compared to AT, JEMs exhibit a significantly greater overlap between the generated and real data distributions. This closer alignment indicates that JEMs more accurately capture the true data distribution, thereby enhancing both generative quality and classification accuracy. Based on these empirical evidences, we derive an key observation: a generative classifier that explicitly minimizes the energy divergence among clean, generated, and adversarial data distribution can effectively bridge the gaps in classification, robustness, and generative capability.
3.3 Energy-based Joint Adversarial Distribution Training
To explicitly minimize the energy divergence among clean, generated, and adversarial data distribution, a straightforward idea is to jointly model the clean data distribution, the adversarial distribution and the classifier via learning the joint probability $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$:
$p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)=p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})p_{\theta}(\tilde{\mathbf{x}},\mathbf{x})=p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})p_{\theta}(\mathbf{x}),$ (6)
where $\theta$ denotes the model parameters. The conditional distribution ${p}_{\theta}\left({y}|\tilde{\mathbf{x}},{\mathbf{x}}\right)$ denotes cross-entropy objective for robust classification, such as AT mkadry2017towards . $p_{\theta}(\tilde{\mathbf{x}},\mathbf{x})$ aims to capture the joint distribution of clean and adversarial examples, and it can be further decomposed using Bayesian theorem as $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})p_{\theta}(\mathbf{x})$. Similar to Eq. 1, the adversarial distribution $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$ can also be parameterized through an energy-based formulation:
$p_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})=\frac{\exp\left(-E_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})\right)}{\tilde{Z}_{\theta}}=\frac{\sum_{y\in\mathbf{y}}\exp\left(f_{\theta}(\tilde{\mathbf{x}})[y]\right)}{\tilde{Z}_{\theta}},$ (7)
where $\tilde{Z}(\theta)=\int_{\tilde{\mathbf{x}}}\exp(-E_{\theta}(\tilde{\mathbf{x}}))$. A natural choice to optimize the model parameters $\theta$ is to maximize the log-likelihood of $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$:
$log\ p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)=log\ p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})+log\ p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})+log\ p_{\theta}(\mathbf{x}),$ (8)
where $log\ p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$ is simply a cross-entropy objective for robust classifier, $log\ p_{\theta}(\mathbf{x})$ can be estimated via various sampling methods welling2011bayesian . In our method, Eq. 2 is approximated by nijkamp2020anatomy :
$\frac{\partial{log}p_{\theta}(\mathbf{x})}{\partial\theta}\approx\frac{\partial}{\partial\theta}\left[\frac{1}{L_{1}}\sum_{i=1}^{L_{1}}E_{\theta}\left(\mathbf{x}_{i}^{+}\right)\right.\\ \left.-\frac{1}{L_{2}}\sum_{i=1}^{L_{2}}E_{\theta}\left(\mathbf{x}_{i}^{-}\right)\right],$ (9)
where $\left\{\mathbf{x}_{i}^{+}\right\}_{i=1}^{L_{1}}$ denote all the training samples in a batch, and $\left\{\mathbf{x}_{i}^{-}\right\}_{i=1}^{L_{2}}$ are i.i.d samples drawn from $p_{\theta}(\mathbf{x})$ via Stochastic Gradient Langevin Dynamics (SGLD) welling2011bayesian :
$\mathbf{x}_{t+1}^{-}=\mathbf{x}_{t}^{-}+\frac{c^{2}}{2}\frac{\partial\log p_{\theta}\left(\mathbf{x}_{t}^{-}\right)}{\partial\mathbf{x}_{t}^{-}}+c\epsilon,c>0,\epsilon\in\mathbf{N}(0,\mathbf{I}),$ (10)
where $c$ is a step size, $\mathbf{N}$ denotes the normal distribution, and $\mathbf{I}$ is the identity matrix. Therefore, the key to optimizing Eq. 8 is to estimate $log\ p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$. Although the full adversarial distribution is not observed during training, we start from a straightforward yet key observation from an energy view: adversarial samples typically appear near the boundaries or even outside the support of the original data distribution, situating them in low-density, thus high-energy regions. Additionally, by visualizing the energy distributions of clean and adversarial samples in Fig. 1 and Tab. 1, we observe that, under robust classifiers, these distributions are nearly indistinguishable. This suggests that, after AT, adversarial examples are pulled back from high-energy area into low-energy (i.e., high-density) areas, close to clean samples. This inspires us to propose a min-max optimization framework to compute $log\ p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$:
$\min_{\theta}\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\max_{\tilde{\mathbf{x}}}\left(E_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})-E_{\theta}(\mathbf{x})\right)\right]\;\text{s.t.}\;\tilde{\mathbf{x}}\;\text{is natural}\;,$ (11)
where $\mathcal{D}$ denotes training data. The inner maximization problem aims to move the adversarial examples out of the high-density area, which corresponds to minimize the joint probability $\log{p_{\theta}}\left(\tilde{\mathbf{x}},y\right)$. To ensure that adversarial examples look more natural, we employ perceptual loss to maintain visual similarity to the clean samples. Specifically, we update SGLD process along the opposite direction of $\Delta_{\mathbf{x}}\log{p_{\theta}}\left(\tilde{\mathbf{x}},y\right)$ to sample the adversarial examples:
$\tilde{\mathbf{x}}_{t+1}=\tilde{\mathbf{x}}_{t}-\frac{c^{2}}{2}\frac{\partial}{\partial\tilde{\mathbf{x}}_{t}}\left[\log{p_{\theta}}\left(\tilde{\mathbf{x}}_{t},y\right)+\lambda d\left(\mathbf{x}_{t},\tilde{\mathbf{x}}_{t}\right)\right],$ (12)
where $\lambda$ is the weight, $d\left(\mathbf{x}_{t},\tilde{\mathbf{x}}_{t}\right)$ denotes the perceptual distance between adversarial and clean samples, which is defined via LPIPSzhang2018unreasonable , which employs a pre-trained network $h(\cdot)$ to compute the feature distances. We begin by extracting activations from $L$ layers of the function $h(\cdot)$ and normalize each across the channel dimension, yielding $\tilde{h}_{l}(x)$ for each layer $l$. Subsequently, each $\tilde{h}_{l}(x)$ is further normalized according to the layer’s size. Using these normalized activations, the distance $d\left(\mathbf{x},\tilde{\mathbf{x}}\right)$ is then defined as follows:
$d(\mathbf{x},\tilde{\mathbf{x}})=\sum_{l}^{L}w_{l}\left\|\frac{\tilde{h}_{l}(\mathbf{x})}{\sqrt{W_{l}H_{l}}}-\frac{\tilde{h}_{l}(\tilde{\mathbf{x}})}{\sqrt{W_{l}H_{l}}}\right\|_{2}^{2},$ (13)
where $W_{l}$ and $H_{l}$ denote the width and height of the activations in layer $l$, $w_{l}$ denotes the weight.
On the other hand, the goal of the outer minimization problem is to find model parameters so that the energy difference between adversarial and clean samples is minimized, i.e. pulling back the adversarial examples back to the high-density area. Unlike existing robust classifiers only maximize the conditional probability $log\ p_{\theta}(y\mid\tilde{\mathbf{x}})$, we capture the full adversarial data distribution during training. Similar for Eq. 9, $\frac{\partial{log}p_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})}{\partial\theta}$ can be approximitated as:
$\frac{\partial{log}p_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})}{\partial\theta}\approx\frac{\partial}{\partial\theta}\left[\frac{1}{L_{1}}\sum_{i=1}^{L_{1}}E_{\theta}\left(\mathbf{x}_{i}^{+}\right)-\frac{1}{L_{2}}\sum_{i=1}^{L_{2}}E_{\theta}\left(\tilde{\mathbf{x}}_{i}|\mathbf{x}_{i}^{+}\right)\right].$ (14)
Our full Algorithm is presented in Algorithm 1.
1 Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M$: sampling iterations; $M_{adv}$: adversarial sampling iterations; $M_{\theta}$: sampling iterations for $\theta$ ;
2 Output: $\{\theta_{1},\dots,\theta_{N}\}$: network weights;
3 Init: randomly initialize $\{\theta_{1},\dots,\theta_{N}\}$;
4
5for $i=1$ to $N_{tra}$ do
6       Sample a mini-batch $\{\mathbf{x},y\}_{i}$;
7       for $m=1$ to $M$ do
8             Obtain ${\mathbf{x}}_{0}$ via random noise;
9             Sample ${\mathbf{x}}_{t}$ via Eq. 10;
10
11       end for
12      Compute $h_{1}=\frac{\partial logp_{\theta}({\mathbf{x}})}{\partial\theta}$ via Eq. 9;
13       //MIN-MAX Optimization for $p_{\theta}(\tilde{{\mathbf{x}}}|{\mathbf{x}})$
14       for $m_{adv}=1$ to $M_{adv}$ do
15             Obtain $\tilde{\mathbf{x}}_{0}$ from ${\mathbf{x}}$ with a perturbation;
16             Sample $\tilde{{\mathbf{x}}}_{t}$ via Eq. 12;
17
18       end for
19      Compute $h_{2}=\frac{\partial\log p_{\theta}(\tilde{{\mathbf{x}}}|{\mathbf{x}})}{\partial\theta}$ via from Eq. 14;
20       Compute $h_{3}=\frac{\partial\log p_{\theta}(y|{\mathbf{x}},\tilde{{\mathbf{x}}})}{\partial\theta}$ via cross-entropy loss;
21       ${h}_{\theta}=h_{1}+h_{2}+h_{3}$;
22       for t = 1 to $M_{\theta}$ do
23             Updata $\theta$;
24
25       end for
26
27 end for
28
29return $\theta$;
Algorithm 1 Energy-based Joint Distribution AT
4 Experiments
4.1 Experimental Setup
We conduct analysis and comparison experiments on several standard classifier datasets, including CIFAR-10 krizhevsky2009learning , CIFAR-100 krizhevsky2009learning . Given that most AT methods utilize WideResNetzagoruyko2016wide architectures, we select the commonly used WRN28-10 as the baseline network. During the entire training process, the learning rate is set to 0.01 and the number of sampling steps is 10. The initialization perturbation for adversarial samples is set to 8/255, and the number of adversarial sampling steps is 5. More details are shown in Appendix B. Note that EB-JDAT is a general optimization framework for JEM that is compatible with various JEM-based modelsgrathwohl2019your ; yang2021jem++ ; yang2023towards ; yang2023m . Therefore, we integrate EB-JDAT with SADAJEMyang2023towards and JEM++yang2021jem++ , two faster and more stable JEMs, to achieve improved performance.
4.2 Comparison Experiments and Analysis
Compared with AT methods
We compare our approach with a range of AT methods, including five popular ones: PGD-AT mkadry2017towards , TRADES zhang2019theoretically , MART wang2019improving , AWP wu2020adversarial , and LBGAT cui2021learnable , as well as four SOTA methods: LAS-AWP jia2022adversarial , TDAT tong2024taxonomy , AGR-TRADES tong2024balancing , and VART zhao2024delving . All AT methods use the WRN34-10 network. For evaluation, we select the PGD mkadry2017towards and AutoAttack (AA) croce2020reliable . The results on CIFAR-10 and CIFAR-100 are shown in Tab. 2, where EB-JDAT not only maintains the classification accuracy of clean data(92.28% on CIFAR-10; 70.19% on CIFAR-100) but also achieves robustness that even surpasses existing SOTA AT methods(+5.96% on CIFAR-10; +2.02% on CIFAR-100). Thus, our method effectively balances the trade-off between accuracy on clean data and robustness against adversarial attacks. Furthermore, EB-JDAT is compatible with both SADAJEMyang2023towards and JEM++yang2021jem++ , thereby demonstrating the universality of our approach.
Table 2: Camparative robustness (%) analysis with different AT methods.
CIFAR-10 CIFAR-100
Method Clean PGD-20 AA Clean PGD-20 AA
Standard Training 96.10 0 0 80.73 0 0
PGD-AT 85.17 55.08 51.69 60.93 26.05 24.33
TRADES 85.72 56.10 53.40 57.10 26.96 24.55
MART 85.17 58.56 51.10 57.29 26.03 23.92
AWP 85.57 58.13 53.90 60.38 33.86 28.86
LBGAT 88.22 54.66 52.23 60.64 34.75 29.33
LAS-AWP 87.74 60.16 55.52 64.89 36.36 30.77
TDAT 82.25 56.03 54.06 57.32 33.17 26.61
AGR-TRADES 86.50 52.56 52.56 58.17 28.34 25.84
VART 84.21 58.32 54.24 / / /
EB-JDAT-JEM++ 91.43 65.69 64.34 70.00 37.14 35.21
EB-JDAT-SADAJEM 92.28 66.12 64.76 70.19 38.38 35.63
Comparisons with AT Using Generated Data.
Another strategy for improving robustness using generative models is to incorporate additional training data generated by generative models in AT pang2022robustness ; gowal2021improving . Therefore we compare EB-JDAT with SCORE pang2022robustness , Better DM wang2023better and the method proposed in gowal2021improving . SOCRE and Better DM utilize the SOTA Diffusion models ho2020denoising to generated addition training data, while gowal2021improving generate addition data from a ensemble of DDPM ho2020denoising , GANs goodfellow2014generative and VDVAE childvery generators. All methods are implemented using the WRN28-10 zagoruyko2016wide . The results are presented in Tab. 3. Although EB-JDAT does not rely on additional generated data or increased training epochs, it still achieves competitive performance in terms of clean accuracy and robustness. EB-JDAT even outperforms SCORE-1M, Better DM-1M and gowal2021improving . Unlike discriminative learning approaches pang2022robustness ; gowal2021improving ; wang2023better , which struggles to capture the underlying structure of both clean and adversarial data distributions, EB-JDAT, as a hybrid generative model, provides more accurate gradient estimations of clean/adversarial data densities, leading to substantial improvements in both accuracy and robustness. Notably, this is achieved without requiring additional training data and with minimal computational overhead.
Table 3: Camparative robustness (%) analysis with AT methods that utilize generative models for data augmentation, P denotes EB-JDAT with perceptual loss.
      Data       Generate data       Epoch       Clean       AA
      CIFAR-10
      SCORE pang2022robustness       1M       400       88.10       61.51
      The method proposed by gowal2021improving       100M       2000       87.50       63.38
      Better DM wang2023better       1M       400       91.12       63.35
      Better DM wang2023better       50M       1600       92.27       67.17
      EB-JDAT-SADAJEM(P)       n/a       100       92.28       64.76
      EB-JDAT-SADAJEM       n/a       100       89.79       66.30
      CIFAR-100
      SCORE pang2022robustness       1M       400       62.08       31.40
      Better DM wang2023better       1M       400       68.06       35.65
      Better DM wang2023better       50M       1600       72.58       38.83
      EB-JDAT-SADAJEM(P)       n/a       100       70.19       35.63
      EB-JDAT-SADAJEM       n/a       100       68.35       38.46
Compared with JEMs and Energy-based AT.
In this section, we compare EB-JDAT with other JEMs grathwohl2019your ; yang2021jem++ ; yang2023towards and energy-based AT zhu2021towards ; mirza2024shedding in terms of classification accuracy, robustness and generative performance. Sec. 4.2 presents an evaluation of our method against JEM variants with respect to classification accuracy and generation quality. To assess generation quality, we employ the Fréchet Inception Distance (FID) and Inception Score (IS) as metrics. Compared to JEMs, EB-JDAT maintains comparable accuracy and generative performance, while EB-JDAT-SADAJEM even outperforms JEM(FID-10.98) and JEM++(FID-9.7) in terms of generation. Moreover, EB-JDAT exhibits significantly improved robustness over JEMs, as shown in Fig. 2. Thus, EB-JDAT bridges the gap between classification, robustness, and generation.
Table 4: Performance comparison of different hybrid models on CIFAR-10, IF denotes informative initialization.
[…]
JEM variants
JEM grathwohl2019your 92.90 38.40 8.76
JEM++ yang2021jem++ 93.73 37.12 8.29
SADAJEM yang2023towards 96.03 11.40 8.63
Energy-based AT
JEAT zhu2021towards 85.16 38.24 8.80
WEAT mirza2024shedding (PCA) 83.36 30.74 8.97
WEAT mirza2024shedding (IF) 83.36 177.92 3.50
EB-JDAT-JEM++ 91.43 38.89 7.66
EB-JDAT-SADAJEM 92.28 27.42 8.05
Figure 2: Comparative robustness (%) analysis with different energy-based AT.
Next, we compare EB-JDAT with energy-based AT zhu2021towards ; mirza2024shedding . First, we compare the accuracy and robustness with them, which is reported in Fig. 2. The results demonstrate that our method significantly enhances robustness with minimal accuracy degradation, outperforming existing energy-based AT. Moreover, our method yields superior generative performance compared to energy-based methods, which are shown in Sec. 4.2. EB-JDAT surpasses JEAT and WEAT with an FID reduction of 10.82 and 3.32, respectively. Unlike JEAT zhu2021towards directly introduces adversarial samples into the JEMs training process to model the joint distribution ${p}_{\theta}(\tilde{\mathbf{x}},y)$, we model the joint probability $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$ to capture the distribution of clean and adversarial samples more accurately, thus performing better in accuracy, robustness and generative ability. On top of that, WEATmirza2024shedding reinterprets TRADESzhang2019theoretically as an EBM to model the conditional distribution $p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$, which is still a discriminative model. Although mirza2024shedding notes that PCA-based initialization can give WEAT generative capabilities(FID=30.74), but under the same comparison of informative initialization, WEAT shows poor generative performance(FID=177.92). Consequently, EB-JDAT outperforms previous energy-based AT in classification accuracy, robustness, and generative capability.
Visual Quality Analysis.
To further visually explore generative ability of EB-JDAT, we generate images via SGLD sampling with informative initialization for conventional AT-trained models, energy-based AT-models, JEMs, and EB-JDAT in Fig. 3. As demonstrated in Fig. 3, the conventional AT-trained models(a, b) and energy-based AT-models(e) possess limited generative capacity, resulting in difficulties in generating high-quality images. Conversely, EB-JDAT produce higher-resolution images with richer background details, which is more closely matching the features of the original dataset. In addition, EB-JDAT(g, h) surpasses JEM(f) and perform competitively with JEM++ and SADAJEM in generative capacity, producing more diverse and detailed images with only a slight reduction in sharpness, particularly in the ’car’ category.
Figure 3: Performance comparison of generated samples from different methods, all methods are sampled by SGLD with informative initialization.
4.3 Ablation Study
Table 5: Ablation study of EB-JDAT-JEM++, ECO denotes the epoch at which collapse occurs.
[…]
1 0 0.1 0.9 90.61 66.41 64.71 146.02 4.44 n/a
2 0.5 0 0.5 74.16 46.28 43.56 69.37 5.95 20
3 0.9 0.1 0 12.59 11.59 11.48 74.29 5.17 44
4 0.35 0.25 0.35 87.13 60.31 58.20 48.52 6.56 70
5 0.45 0.05 0.45 91.43 65.69 64.34 38.89 7.66 n/a
In EB-JDAT, we employ a weighted combination of gradients, as detailed in Algorithm 1, to compute the overall optimization gradient, expressed as $h_{\theta}=w_{1}h_{1}+w_{2}h_{2}+w_{3}h_{3}$. To assess the individual contribution of each weight, we perform ablation studies by setting each of $w_{1}$, $w_{2}$, and $w_{3}$ to zero and evaluating their impact on overall performance. The results, presented in Table 5, reveal two key findings: (1) When $w_{2}=0$, the optimization focuses solely on classification loss and the clean-sample distribution. Although this setting enhances robustness, the significant energy gap between clean and adversarial samples prevents model convergence, leading to model collapse by the 20th epoch. (2) Maintaining a small value for $w_{2}$ not only balances accuracy and robustness but also stabilizes the training phase, reducing the risk of model collapse. Consequently, experimental results indicate that $w_{1}=0.45$, $w_{2}=0.05$, and $w_{3}=0.45$ yields optimal performance in terms of accuracy, robustness, generative capability, and stability. This configuration is therefore adopted as the default setting.
5 Limitation And Broader Impact
Training EB-JDAT on complex, high-dimensional data remains challenging. This challenge is also encountered by JEMs,including JEM grathwohl2019your , JEM++ yang2021jem++ and SADAJEM yang2023towards . This instability arises from the typically sharp probability distribution of real data in high-dimensional space, which leads to inaccurate guidance for image sampling in regions with low data density. Although the approach we propose is a general and flexible optimization framework for all JEMs, considering training stability, we recommend training within faster and more stable JEM variants, such as JEM++ and SADAJEM. Nevertheless, our approach allows a standard classifier to achieve good generative ability and robustness without sacrificing much accuracy, resulting in the best overall performance in terms of accuracy, generative capability, and robustness.
From a societal perspective, a robustness generative-classifier more effectively mitigates data scarcity and adversarial perturbations. However, if the training dataset itself is high-risk, the model may generate equally high-risk content, potentially posing harm to society. To minimize such harm, we will soon release user guidelines and begin implementing safety filters.
6 Conclusion
In this work, we systematically investigate the underlying causes of the performance gap between discriminate classifiers, robust classifiersmkadry2017towards ; zhang2019theoretically and JEM-based classifiers grathwohl2019your ; yang2023towards in terms of accuracy, generative capability, and adversarial robustness, and propose a general and flexible optimization framework, called EB-JDAT to eliminate the inherent trilemma. Our experiments validate the effectiveness of these techniques across various benchmarks, demonstrating superior performance in achieving the triple trade-off, with particularly impressive SOTA results in robustness.
Appendix A Ethics Statement
We exclusively employed publicly available datasets and models for research purposes and did not involve human subjects. EB-JDAT is a robust generative discriminator with powerful image-generation capabilities; if misused, it could produce content that poses societal risks, and no safeguards currently limit its outputs. Nevertheless, our approach achieves significantly greater robustness than earlier generative discriminators, ensuring its safe deployment in real-world applications. We therefore contend that the benefits of our work far outweigh any potential harms arising from disclosure of our methods. In addition, a fundamental classifier endowed with generative capacity and discriminative robustness more effectively mitigates data scarcity and adversarial perturbations.
Appendix B Experimental details
The experiments were carried out on a platform featuring an AMD EPYC 7542 32-core CPU with a clock speed of 2039.813 GHz and an NVIDIA GeForce RTX 3090 GPU, equipped with 24 GB of RAM. The proposed method was implemented using the open-source PyTorch machine learning framework. In addition, we use the same hyperparameter settings and validation-set configuration for training and evaluation as in JEM variants [2, 3].

### v2

@@
 Abstract
-Joint Energy-based Models (JEMs), a class of hybrid generative-discriminative models, are well known for their ability to achieve both high classification accuracy and generative capability within a single model.
-However, their robustness still lags significantly behind the classifiers based adversarial training (AT).
-Conversely, while AT is currently the most effective approach to improving the classifier’s robustness, it typically sacrifices accuracy on clean data and lacks generative capability.
-The triple trade-off between classification accuracy, generative capability and robustness, raises a natural question: Can a single model simultaneously achieve high classification accuracy, adversarial robustness, and generative performance?
-– a goal that has been rarely explored.
-To address this question, we systematically analyze the energy distribution differences of clean, adversarial, and generated samples across various JEM variants and adversarially trained models.
-We observe that AT tends to reduce the energy gap between clean and adversarial samples, while JEMs reduce the gap between clean and synthetic ones.
-This observation suggests a key insight: if the energy distributions of all three data types can be aligned, we might unify the strengths of AT and JEMs, resolving their inherent trade-offs.
-Building on this idea, we propose Energy-based Joint Distribution Adversarial Training (EB-JDAT), to jointly model the clean data distribution, the adversarial distribution, and the classifier by maximizing their joint probability.
-To capture the full adversarial distribution, we introduce a new min-max energy optimization strategy.
-EB-JDAT is a general and flexible optimization method, compatible with various JEM variants.
-Extensive experimental results demonstrate that EB-JDAT not only maintains near original accuracy and generative capability of JEMs, but also significantly enhances robustness, even surpassing state-of-the-art ATs.
+Joint Energy-based Models (JEMs) are well known for their ability to unify classification and generation within a single framework.
+Despite their promising generative and discriminative performance, their robustness remains far inferior to adversarial training (AT), which, conversely, achieves strong robustness but sacrifices clean accuracy and lacks generative ability.
+This inherent trilemma—balancing classification accuracy, robustness, and generative capability—raises a fundamental question: Can a single model achieve all three simultaneously?
+To answer this, we conduct a systematic energy landscape analysis of clean, adversarial, and generated samples across various JEM and AT variants.
+We observe that AT reduces the energy gap between clean and adversarial samples, while JEMs narrow the gap between clean and synthetic ones.
+This observation suggests a key insight: if the energy distributions of all three data types can be aligned, we might bridge their performance disparities.
+Building on this idea, we propose Energy-based Joint Distribution Adversarial Training (EB-JDAT), a unified generative-discriminative-robust framework that maximizes the joint probability of clean and adversarial distribution.
+EB-JDAT introduces a novel min–max energy optimization to explicitly aligning energies across clean, adversarial, and generated samples.
+Extensive experiments on CIFAR-10, CIFAR-100, and ImageNet subsets demonstrate that EB-JDAT achieves state-of-the-art robustness while maintaining near-original accuracy and generation quality of JEMs, effectively resolving the triple trade-off between accuracy, robustness, and generation.
 1 Introduction
-Recently, Joint Energy-based Models (JEMs) grathwohl2019your ; yang2021jem++ ; yang2023towards have garnered significant attention for their ability to integrate generative modeling within a discriminative framework.
+Figure 1: Comparisons of SOTA AT-based methods on CIFAR-10 in terms of accuracy and robustness (AutoAttack). Our method achieves the best robustness while maintaining competitive standard accuracy.
+Recently, Joint Energy-based Models (JEMs) [13, 39, 41] have garnered significant attention for their ability to integrate generative modeling within a discriminative framework.
 This approach effectively bridges classification and generation from an energy-based perspective and exhibits surprising inherent robustness.
-However, JEMs still exhibit limited robustness compared to Adversarial Training (AT) mkadry2017towards .
-On the other hand, AT jia2022adversarial ; hammoudeh2024provable ; bortolussi2024robustness ; fang20253sat is widely recognized as the most effective method for improving the robustness of discriminative classifiers.
-However, it typically sacrifices accuracy on clean data and lacks generative capability zhang2019theoretically .
+However, JEMs still exhibit limited robustness compared to Adversarial Training (AT) [25].
+On the other hand, AT [19, 14, 2, 45] is widely recognized as the most effective method for improving the robustness of discriminative classifiers.
+However, it typically sacrifices accuracy on clean data and lacks generative capability [44].
 Consequently, the inherent gap in classification, robustness and generation raises a natural question: Can a single model simultaneously achieve high classification accuracy, adversarial robustness, and generative performance?
@@
 indicates generated samples, indicates adversarial samples, indicates clean samples.
-Table 1: Mean and variance of one-to-one energy differences between clean and adv.
+To answer this question, we investigate the underlying causes of the performance gap between AT [25, 44] and JEMs [13, 41] in terms of accuracy, generative capability, and adversarial robustness from an energy perspective.
+We compute the energy distributions of clean, adversarial, and generated samples, with results shown in Fig.
+2.
+Further, we compute the mean and variance of the one-to-one energy differences between clean and adversarial samples, as reported in Tab.
+1.
+The energy distributions shows that AT reduces the energy gap between clean and adversarial samples, thereby bringing robustness.
+In contrast, JEMs reduce the gap between clean and generated samples, which brings generative capability and higher accuracy.
+Based on these insights, we draw a key conclusion: if the energy distributions of all three data types can be aligned, we might unify the strengths of AT and JEMs, resolving their inherent trilemma.
+Table 1: Mean and variance of one‐to‐one energy differences between clean and adv. Results in bold indicate the best.
 […]
 Standard 10.18 3.27 0.06
-PGD-AT 1.46 0.83 55.08
-TRADES 1.01 0.45 56.10
-JEM 5.93 7.30 7.28
-SADAJEM 2.66 3.83 42.46
+PGD-AT [25] 1.46 0.83 55.08
+TRADES [44] 1.01 0.45 56.10
+JEM [13] 5.93 7.30 7.28
+SADAJEM [41] 2.66 3.83 43.64
 EB-JDAT 0.89 0.31 66.12
-To answer this question, we begin by investigating, from an energy perspective, the underlying causes of the performance gap between AT mkadry2017towards ; zhang2019theoretically and JEMs grathwohl2019your ; yang2023towards in terms of accuracy, generative capability, and adversarial robustness.
-We compute the energy distributions of clean vs.
-adversarial vs.
-generated samples by utilizing different models, and the results are shown in Fig.
-1.
-First, the energy distributions of clean vs.
-adversarial samples, give a clear indication of AT tends to reduce the energy gap between clean and adversarial samples.
-Interestingly, JEMs implicitly reduce the gap between them, even though the distribution of adversarial samples is not explicitly incorporated during training, which explains the potential source of robustness in JEMs.
-Further, we compute the mean and variance of one-to-one energy differences between them.
-As reported in Tab.
-1, models with smaller energy differences exhibit higher robustness.
-Next, we investigate the classification and generation gaps by analyzing the energy distributions of clean vs.
-generated samples in Fig.
-1.
-Compared to AT, JEMs reduce the gap between them, achieving high accuracy in classification while maintaining strong generative quality.
-Building on these observations, we draw a key insight: if the energy distributions of all three data types can be aligned, we might unify the strengths of AT and JEMs, resolving their inherent trilemma.
-Inspired by this idea, we align the energy distribution of the three data types by jointly modeling the clean data distribution, the adversarial distribution and the classifier, by learning $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$ via energy-based model, where $\mathbf{x}$, $y$ and $\tilde{\mathbf{x}}$ denote the clean samples, labels and adversarial examples, and $\theta$ denote the model parameters.
+Inspired by this idea, we align the energy distributions of clean, adversarial and generated data through a joint energy-based model of $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$, with $\mathbf{x}$ clean samples, $\tilde{\mathbf{x}}$ adversarial examples, $y$ class labels and $\theta$ the model parameters.
 Through Bayesian decomposition, $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$ can be further factorized as the data distribution $p_{\theta}(\mathbf{x})$, adversarial distribution $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$ and $p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$.
-While $p_{\theta}(\mathbf{x})$ can be computed via approximated sampling methods nijkamp2020anatomy and $p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$ is simply a cross-entropy objective for robust classification, computing $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$ is not straightforward because the full adversarial distribution is not observed during the training phase.
-However, our key observation is that adversarial perturbations almost surely move samples out of the original data manifold, which lies in high-density areas, into low-density areas where the classifier is more likely to fail.
+While $p_{\theta}(\mathbf{x})$ can be computed via approximated sampling methods [27] and $p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$ is simply a cross-entropy objective for robust classification, computing $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$ is not straightforward because the full adversarial distribution is not observed during the training phase.
+However, our key observation is that adversarial perturbations almost surely move samples out of the original data manifold, which lies in high-density areas, into low-density areas where the classifier is more likely to fail, as illustrated by the weakly robust classifiers JEM [13] and SADAJEM [41] in Fig.
+2.
 Motivated by this observation, we propose a new min-max energy optimization to approximate $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$.
@@
 We name our proposed method as Energy-based Joint Distribution Adversarial Training, or EB-JDAT.
-Extensive experiments across multiple datasets and attacks show that EB-JDAT achieves 64.76% and 35.63% robustness (under AutoAttack) on CIFAR-10 and CIFAR-100 datasets, outperforming the SOTA AT methods by +9.24% and +4.86% respectively, while maintaining near original clean accuracy and generative capability, which is competitive to most advanced JEMs.
+Extensive experiments across multiple datasets and attacks show that EB-JDAT achieves 68.76%, 35.63% and 32.40% robustness (under AutoAttack [6]) on CIFAR-10, CIFAR-100 and ImageNet subset, outperforming the SOTA AT methods by +10.78%, +4.70% and +7.88% respectively, while maintaining near original clean accuracy and generative capability, which is competitive to most advanced JEMs.
 2 Related Work
 Adversarial Defense.
-Since the discovery of adversarial vulnerabilities in deep learning models yuan2021automa ; ma2023transferable , numerous defense strategies have been proposed to mitigate this security challenge.
-Prominent approaches include input denoising li2022feature , gradient regularization li2023revisiting ; jia2022prior , defensive distillation huang2023boosting , and adversarial training (AT) mkadry2017towards ; fang20253sat .
-Recent studies confirm that AT remains among the most effective defense methods sengphanith2023evaluating , significantly enhancing model robustness by incorporating adversarial examples into the training process.
-Notably, mkadry2017towards first formulated AT as a saddle-point optimization problem (PGD), laying the foundational framework for subsequent robust training methodologies.
-As adversarial attack methods have evolved, various novel strategies have emerged, refining conventional AT methods jia2022adversarial ; hammoudeh2024provable ; bortolussi2024robustness ; fang20253sat .
+Since the discovery of adversarial vulnerabilities in deep learning models [42, 24], numerous defense strategies have been proposed to mitigate this security challenge.
+Prominent approaches include input denoising [21], gradient regularization [22, 18], defensive distillation [16], and adversarial training (AT) [25, 10, 45].
+Recent studies confirm that AT remains among the most effective defense methods [29], significantly enhancing model robustness by incorporating adversarial examples into the training process.
+Notably, [25] first formulated AT as a saddle-point optimization problem (PGD), laying the foundational framework for subsequent robust training methodologies.
+As adversarial attack methods have evolved, various novel strategies have emerged, refining conventional AT methods [19, 14, 2, 10, 45].
 However, many of these strategies prioritize robustness at the expense of clean accuracy.
 which is the principal evaluation metric in visual classification, highlighting an inherent trade-off in AT.
-To mitigate this issue, zhang2019theoretically proposed TRADES, which constrains the divergence in output probability between clean and adversarial examples.
-Additionally, co2022jacobian combined Jacobian regularization with model ensembling, introducing the Jacobians ensemble approach that effectively balances robustness and accuracy against universal adversarial perturbations.
-Furthermore, DUCAT wang2024new assigns a dummy class to each original class to absorb hard adversarial examples during training and remaps them back at inference to mitigating the trade-off.
+To mitigate this issue, [44] proposed TRADES, which constrains the divergence in output probability between clean and adversarial examples.
+Additionally, [4] combined Jacobian regularization with model ensembling, introducing the Jacobians ensemble approach that effectively balances robustness and accuracy against universal adversarial perturbations.
+Furthermore, DUCAT [32] assigns a dummy class to each original class to absorb hard adversarial examples during training and remaps them back at inference to mitigating the trade-off.
 Although these approaches represent notable advancements, they still suffer from substantial accuracy degradation relative to standard classifiers, while offering only modest enhancements in robustness.
@@
 Robust Energy-based Model.
-Recent studies have indicated that features extracted by robust classifiers are more closely aligned with human perceptual mechanisms compared to those derived from standard classifiers wangunified .
-The Joint Energy-based Model (JEM) grathwohl2019your reformulates conventional softmax classifiers into an energy-based model (EBM), facilitating hybrid discriminative-generative modeling and clarifying the connection between robustness and generative modeling.
-More recent efforts have focused on improving training efficiency and speed in JEM yang2021jem++ ; yang2023towards , effectively bridging the gap between classification and generation.
+Recent studies have indicated that features extracted by robust classifiers are more closely aligned with human perceptual mechanisms compared to those derived from standard classifiers [33].
+The Joint Energy-based Model (JEM) [13] reformulates conventional softmax classifiers into an energy-based model (EBM), facilitating hybrid discriminative-generative modeling and clarifying the connection between robustness and generative modeling.
+More recent efforts have focused on improving training efficiency and speed in JEM [39, 41], effectively bridging the gap between classification and generation.
 Additionally, several approaches specifically focus on enhancing robustness, from an energy perspective.
-JEAT zhu2021towards initially demonstrated a link between AT and EBM, highlighting that despite differences in how these methods modify the energy function, they ultimately employ comparable contrastive strategies.
-Subsequently, mirza2024shedding reinterpreted robust classifiers within an energy-based framework and analyzed the energy landscape to reveal differential effects of targeted and non-targeted attacks during AT.
+JEAT [46] initially demonstrated a link between AT and EBM, highlighting that despite differences in how these methods modify the energy function, they ultimately employ comparable contrastive strategies.
+Subsequently, [26] reinterpreted robust classifiers within an energy-based framework and analyzed the energy landscape to reveal differential effects of targeted and non-targeted attacks during AT.
 However, these methods often sacrifice robustness in exchange for notable reductions in accuracy and generative capacity, with some even proving less robust than traditional adversarial defense strategies.
@@
 3.1 Preliminaries
-Energy-based Models (EBMs)
-From an energy perspective, the clean data distribution ${p}_{\theta}(\mathbf{x})$ can be explicitly parameterized through an energy function grathwohl2019your :
+Energy-based Models (EBMs).
+From an energy perspective, the clean data distribution ${p}_{\theta}(\mathbf{x})$ can be explicitly parameterized through an energy function [13]:
 $p_{\theta}(\mathbf{x})=\frac{\exp\left(-E_{\theta}(\mathbf{x})\right)}{Z(\theta)}=\frac{\sum_{y\in\mathbf{y}}\exp\left(f_{\theta}(\mathbf{x})[y]\right)}{Z(\theta)},$ (1)
-where $E_{\theta}(\cdot):\mathbb{R}^{D}\rightarrow\mathbb{R}$ denotes the energy function.
+where $E_{\theta}(\cdot):\mathbb{R}^{D}\rightarrow\mathbb{R}$ denotes the energy function, $f_{\theta}(x)[y]$ denotes the $y^{\mathrm{th}}$ index of $f_{\theta}[y]$, i.e., the logit corresponding the $y^{\mathrm{th}}$ class label.
 This energy function assigns each data point a scalar value representing its density within the continuous data distribution.
@@
 To optimize the energy-based model, a common strategy is to perform maximum likelihood estimation of the parameters $\theta$, in which the gradient of the log-likelihood is expressed as follows:
-$\frac{\partial{log}p_{\theta}(\mathbf{x})}{\partial\theta}=\mathbb{E}_{p_{\theta}(\mathbf{x^{\prime}})}\left[\frac{\partial E_{\theta}(\mathbf{x^{\prime}})}{\partial\theta}\right]-\frac{\partial E_{\theta}(\mathbf{x})}{\partial\theta},$ (2)
-where $\mathbf{x^{\prime}}$ is simply a sample obtained by sampling from the model distribution.
+$\frac{\log p_{\theta}(\mathbf{x})}{\partial\theta}=\mathbb{E}_{p_{\theta}(\mathbf{x^{\prime}})}\left[\frac{\partial E_{\theta}(\mathbf{x^{\prime}})}{\partial\theta}\right]-\frac{\partial E_{\theta}(\mathbf{x})}{\partial\theta},$ (2)
+where $\mathbf{x^{\prime}}$ is simply a sample obtained by sampling the distribution of models.
 Joint Energy-based Models (JEMs)
-JEM grathwohl2019your reinterprets the standard softmax classifier within the framework of an EBM.
+JEM [13] reinterprets the standard softmax classifier within the framework of an EBM.
 Specifically, it redefines the logits of the classifier $f_{\theta}$ to model the joint distribution $p_{\theta}(\textbf{x},y)$ of the input data x and their labels $y$ using an energy-based approach:
 $p_{\theta}(\mathbf{x},y)=\frac{\exp\left(f_{\theta}(\mathbf{x})[y]\right)}{Z(\theta)}.$ (3)
-In JEMs grathwohl2019your ; yang2021jem++ ; yang2023towards , the model parameters $\theta$ are optimized by maximizing $\log p_{\theta}(\mathbf{x},y)$:
+In JEMs [13, 39, 41], the model parameters $\theta$ are optimized by maximizing $\log p_{\theta}(\mathbf{x},y)$:
 $\log p_{{\theta}}(\mathbf{x},y)=\log p_{{\theta}}(y\mid\mathbf{x})+\log p_{{\theta}}(\mathbf{x}).$ (4)
-Adversarial Training (AT)
-Adversarial training can be viewed as a min-max optimization problem that integrates inner loss maximization and outer loss minimization mkadry2017towards .
+Adversarial Training (AT).
+Adversarial training can be viewed as a min-max optimization problem that integrates inner loss maximization and outer loss minimization [25].
 Given data $\mathbf{x}\in\mathbf{X}$, its corresponding label $y\in\mathbf{y}$, and a classification model $f_{\theta}$, the objective of the inner maximization is to identify the most aggressive adversarial examples by maximizing the loss.
@@
 3.2 Motivation
-ATmkadry2017towards has demonstrated strong robustness against adversarial attacks.
-However, the adversarial samples introduced during training lie outside the true data manifold, causing the model to overemphasize these adversarial features while neglecting meaningful patterns in clean data ilyas2019adversarial .
+AT[25] has demonstrated strong robustness against adversarial attacks.
+However, the adversarial samples introduced during training lie outside the true data manifold, causing the model to overemphasize these adversarial features while neglecting meaningful patterns in clean data [17].
 Consequently, AT-trained models exhibit a noticeable accuracy gap compared to standard classifiers.
-In contrast, JEMs have shown the ability to reinterpret discriminative classifiers as generative models grathwohl2019your , achieving notable performance in both image classification and generation.
+In contrast, JEMs have shown the ability to reinterpret discriminative classifiers as generative models [13], achieving notable performance in both image classification and generation.
 While JEMs exhibit improved adversarial robustness relative to standard models, a significant robustness gap still remains when compared to AT-trained methods.
@@
 generated samples to investigate the source of the performance gains in JEMs and AT, which are shown in Fig.
-1.
+2.
+Analysis of the Robustness.
 We first analyze the energy distributions of clean vs.
@@
 This also explains the source of robustness in JEMs.
+Analysis of the Generative Capability.
 Next, we investigate the classification and generative gaps by analyzing the energy distributions of real vs.
 generated samples in Fig.
-1.
+2.
 Compared to AT, JEMs exhibit a significantly greater overlap between the generated and real data distributions.
@@
 Based on these empirical evidences, we derive an key observation: a generative classifier that explicitly minimizes the energy divergence among clean, generated, and adversarial data distribution can effectively bridge the gaps in classification, robustness, and generative capability.
-3.3 Energy-based Joint Adversarial Distribution Training
+3.3 Energy-based Joint Distribution Adversarial Training
 To explicitly minimize the energy divergence among clean, generated, and adversarial data distribution, a straightforward idea is to jointly model the clean data distribution, the adversarial distribution and the classifier via learning the joint probability $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$:
-$p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)=p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})p_{\theta}(\tilde{\mathbf{x}},\mathbf{x})=p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})p_{\theta}(\mathbf{x}),$ (6)
+$\begin{split}p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)&=p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})\,p_{\theta}(\tilde{\mathbf{x}},\mathbf{x})\\ &=p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})\,p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})\,p_{\theta}(\mathbf{x}),\end{split}$ (6)
 where $\theta$ denotes the model parameters.
-The conditional distribution ${p}_{\theta}\left({y}|\tilde{\mathbf{x}},{\mathbf{x}}\right)$ denotes cross-entropy objective for robust classification, such as AT mkadry2017towards .
+The conditional distribution ${p}_{\theta}\left({y}|\tilde{\mathbf{x}},{\mathbf{x}}\right)$ denotes cross-entropy objective for robust classification, such as AT [25].
 $p_{\theta}(\tilde{\mathbf{x}},\mathbf{x})$ aims to capture the joint distribution of clean and adversarial examples, and it can be further decomposed using Bayesian theorem as $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})p_{\theta}(\mathbf{x})$.
@@
 1, the adversarial distribution $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$ can also be parameterized through an energy-based formulation:
-$p_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})=\frac{\exp\left(-E_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})\right)}{\tilde{Z}_{\theta}}=\frac{\sum_{y\in\mathbf{y}}\exp\left(f_{\theta}(\tilde{\mathbf{x}})[y]\right)}{\tilde{Z}_{\theta}},$ (7)
-where $\tilde{Z}(\theta)=\int_{\tilde{\mathbf{x}}}\exp(-E_{\theta}(\tilde{\mathbf{x}}))$.
+$p_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})=\frac{\exp\left(-E_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})\right)}{\tilde{Z}_{\theta}}=\frac{\sum_{y\in\mathbf{y}}\exp\left(f_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})[y]\right)}{\tilde{Z}_{\theta}},$ (7)
+where $\tilde{Z}(\theta)=\int_{\tilde{\mathbf{x}}}\exp(-E_{\theta}(\tilde{\mathbf{x}}|\mathbf{x}))$.
 A natural choice to optimize the model parameters $\theta$ is to maximize the log-likelihood of $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$:
-$log\ p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)=log\ p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})+log\ p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})+log\ p_{\theta}(\mathbf{x}),$ (8)
-where $log\ p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$ is simply a cross-entropy objective for robust classifier, $log\ p_{\theta}(\mathbf{x})$ can be estimated via various sampling methods welling2011bayesian .
+$\log p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$ $=\log p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$
+$\quad+\log p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})+\log p_{\theta}(\mathbf{x}),$ (8)
+where $\log p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$ is simply a cross-entropy objective for robust classifier, $\log p_{\theta}(\mathbf{x})$ can be estimated via various sampling methods [36].
 In our method, Eq.
-2 is approximated by nijkamp2020anatomy :
-$\frac{\partial{log}p_{\theta}(\mathbf{x})}{\partial\theta}\approx\frac{\partial}{\partial\theta}\left[\frac{1}{L_{1}}\sum_{i=1}^{L_{1}}E_{\theta}\left(\mathbf{x}_{i}^{+}\right)\right.\\ \left.-\frac{1}{L_{2}}\sum_{i=1}^{L_{2}}E_{\theta}\left(\mathbf{x}_{i}^{-}\right)\right],$ (9)
-where $\left\{\mathbf{x}_{i}^{+}\right\}_{i=1}^{L_{1}}$ denote all the training samples in a batch, and $\left\{\mathbf{x}_{i}^{-}\right\}_{i=1}^{L_{2}}$ are i.i.d samples drawn from $p_{\theta}(\mathbf{x})$ via Stochastic Gradient Langevin Dynamics (SGLD) welling2011bayesian :
+2 is approximated by [27]:
+$\frac{\partial{\log}p_{\theta}(\mathbf{x})}{\partial\theta}\approx\frac{\partial}{\partial\theta}\left[\frac{1}{L_{1}}\sum_{i=1}^{L_{1}}E_{\theta}\left(\mathbf{x}_{i}^{+}\right)\right.\\ \left.-\frac{1}{L_{2}}\sum_{i=1}^{L_{2}}E_{\theta}\left(\mathbf{x}_{i}^{-}\right)\right],$ (9)
+where $\left\{\mathbf{x}_{i}^{+}\right\}_{i=1}^{L_{1}}$ denote all the training samples in a batch, and $\left\{\mathbf{x}_{i}^{-}\right\}_{i=1}^{L_{2}}$ are i.i.d samples drawn from $p_{\theta}(\mathbf{x})$ via Stochastic Gradient Langevin Dynamics (SGLD) [36]:
 $\mathbf{x}_{t+1}^{-}=\mathbf{x}_{t}^{-}+\frac{c^{2}}{2}\frac{\partial\log p_{\theta}\left(\mathbf{x}_{t}^{-}\right)}{\partial\mathbf{x}_{t}^{-}}+c\epsilon,c>0,\epsilon\in\mathbf{N}(0,\mathbf{I}),$ (10)
 where $c$ is a step size, $\mathbf{N}$ denotes the normal distribution, and $\mathbf{I}$ is the identity matrix.
-Therefore, the key to optimizing Eq.
-8 is to estimate $log\ p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$.
+Therefore, the key to optimizing Sec.
+3.3 is to estimate $\log\ p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$.
 Although the full adversarial distribution is not observed during training, we start from a straightforward yet key observation from an energy view: adversarial samples typically appear near the boundaries or even outside the support of the original data distribution, situating them in low-density, thus high-energy regions.
@@
 This suggests that, after AT, adversarial examples are pulled back from high-energy area into low-energy (i.e., high-density) areas, close to clean samples.
-This inspires us to propose a min-max optimization framework to compute $log\ p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$:
-$\min_{\theta}\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\max_{\tilde{\mathbf{x}}}\left(E_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})-E_{\theta}(\mathbf{x})\right)\right]\;\text{s.t.}\;\tilde{\mathbf{x}}\;\text{is natural}\;,$ (11)
+This inspires us to propose a min-max optimization framework to compute $\log p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$:
+$\min_{\theta}\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\max_{\|\tilde{\mathbf{x}}-\mathbf{x}\|\in\Omega}\left(E_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})-E_{\theta}(\mathbf{x})\right)\right],$ (11)
 where $\mathcal{D}$ denotes training data.
-The inner maximization problem aims to move the adversarial examples out of the high-density area, which corresponds to minimize the joint probability $\log{p_{\theta}}\left(\tilde{\mathbf{x}},y\right)$.
-To ensure that adversarial examples look more natural, we employ perceptual loss to maintain visual similarity to the clean samples.
-Specifically, we update SGLD process along the opposite direction of $\Delta_{\mathbf{x}}\log{p_{\theta}}\left(\tilde{\mathbf{x}},y\right)$ to sample the adversarial examples:
-$\tilde{\mathbf{x}}_{t+1}=\tilde{\mathbf{x}}_{t}-\frac{c^{2}}{2}\frac{\partial}{\partial\tilde{\mathbf{x}}_{t}}\left[\log{p_{\theta}}\left(\tilde{\mathbf{x}}_{t},y\right)+\lambda d\left(\mathbf{x}_{t},\tilde{\mathbf{x}}_{t}\right)\right],$ (12)
-where $\lambda$ is the weight, $d\left(\mathbf{x}_{t},\tilde{\mathbf{x}}_{t}\right)$ denotes the perceptual distance between adversarial and clean samples, which is defined via LPIPSzhang2018unreasonable , which employs a pre-trained network $h(\cdot)$ to compute the feature distances.
-We begin by extracting activations from $L$ layers of the function $h(\cdot)$ and normalize each across the channel dimension, yielding $\tilde{h}_{l}(x)$ for each layer $l$.
-Subsequently, each $\tilde{h}_{l}(x)$ is further normalized according to the layer’s size.
-Using these normalized activations, the distance $d\left(\mathbf{x},\tilde{\mathbf{x}}\right)$ is then defined as follows:
-$d(\mathbf{x},\tilde{\mathbf{x}})=\sum_{l}^{L}w_{l}\left\|\frac{\tilde{h}_{l}(\mathbf{x})}{\sqrt{W_{l}H_{l}}}-\frac{\tilde{h}_{l}(\tilde{\mathbf{x}})}{\sqrt{W_{l}H_{l}}}\right\|_{2}^{2},$ (13)
-where $W_{l}$ and $H_{l}$ denote the width and height of the activations in layer $l$, $w_{l}$ denotes the weight.
-On the other hand, the goal of the outer minimization problem is to find model parameters so that the energy difference between adversarial and clean samples is minimized, i.e.
-pulling back the adversarial examples back to the high-density area.
-Unlike existing robust classifiers only maximize the conditional probability $log\ p_{\theta}(y\mid\tilde{\mathbf{x}})$, we capture the full adversarial data distribution during training.
-Similar for Eq.
-9, $\frac{\partial{log}p_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})}{\partial\theta}$ can be approximitated as:
-$\frac{\partial{log}p_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})}{\partial\theta}\approx\frac{\partial}{\partial\theta}\left[\frac{1}{L_{1}}\sum_{i=1}^{L_{1}}E_{\theta}\left(\mathbf{x}_{i}^{+}\right)-\frac{1}{L_{2}}\sum_{i=1}^{L_{2}}E_{\theta}\left(\tilde{\mathbf{x}}_{i}|\mathbf{x}_{i}^{+}\right)\right].$ (14)
+The inner maximization problem aims to move the adversarial examples out of the high-density area, which corresponds to minimize the joint probability $\log{p_{\theta}}\left((\tilde{\mathbf{x}}\mid\mathbf{x}),y\right)$.
+Specifically, we update SGLD process along the opposite direction of $\nabla_{\mathbf{x}}\log{p_{\theta}}\left((\tilde{\mathbf{x}}\mid\mathbf{x}),y\right)$ to sample the adversarial examples:
+$\tilde{\mathbf{x}}_{t+1}=\tilde{\mathbf{x}}_{t}-\frac{c^{2}}{2}\frac{\partial\log{p_{\theta}}\left((\tilde{\mathbf{x}}\mid\mathbf{x}),y\right)}{\partial\tilde{\mathbf{x}}_{t}},$ (12)
+[…]
+$\frac{\partial{\log}p_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})}{\partial\theta}\approx\frac{\partial}{\partial\theta}\left[\frac{1}{L_{1}}\sum_{i=1}^{L_{1}}E_{\theta}\left(\mathbf{x}_{i}^{+}\right)-\frac{1}{L_{2}}\sum_{i=1}^{L_{2}}E_{\theta}\left(\tilde{\mathbf{x}}_{i}|\mathbf{x}_{i}^{+}\right)\right].$ (13)
 Our full Algorithm is presented in Algorithm 1.
-1 Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M$: sampling iterations; $M_{adv}$: adversarial sampling iterations; $M_{\theta}$: sampling iterations for $\theta$ ;
-2 Output: $\{\theta_{1},\dots,\theta_{N}\}$: network weights;
-3 Init: randomly initialize $\{\theta_{1},\dots,\theta_{N}\}$;
+1Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M$: sampling iterations; $M_{adv}$: adversarial sampling iterations; $M_{\theta}$: sampling iterations for $\theta$ ;
+2 Output: $\theta$: network weight;
+3 Init: randomly initialize $\theta$;
 4
@@
 6   Sample a mini-batch $\{\mathbf{x},y\}_{i}$;
+7   Obtain ${\mathbf{x}}_{0}$ via random noise;
 8 for $m=1$ to $M$ do
-8             Obtain ${\mathbf{x}}_{0}$ via random noise;
 9      Sample ${\mathbf{x}}_{t}$ via Eq. 10;
@@
 11   end for
-12      Compute $h_{1}=\frac{\partial logp_{\theta}({\mathbf{x}})}{\partial\theta}$ via Eq. 9;
+12  Compute $h_{1}=\frac{\partial\log p_{\theta}({\mathbf{x}})}{\partial\theta}$ via Eq. 9;
 13 //MIN-MAX Optimization for $p_{\theta}(\tilde{{\mathbf{x}}}|{\mathbf{x}})$
+14   Obtain $\tilde{\mathbf{x}}_{0}$ from $\mathbf{x}$} with a perturbation;
 15 for $m_{adv}=1$ to $M_{adv}$ do
-15             Obtain $\tilde{\mathbf{x}}_{0}$ from ${\mathbf{x}}$ with a perturbation;
 16      Sample $\tilde{{\mathbf{x}}}_{t}$ via Eq. 12;
@@
 22 for t = 1 to $M_{\theta}$ do
-23             Updata $\theta$;
-24
+23      update $\theta$ via ${h}_{\theta}$;
 24   end for
-26
+25
 26 end for
-28
-29return $\theta$;
+27
+28return $\theta$;
+29
 Algorithm 1 Energy-based Joint Distribution AT
@@
 4.1 Experimental Setup
-We conduct analysis and comparison experiments on several standard classifier datasets, including CIFAR-10 krizhevsky2009learning , CIFAR-100 krizhevsky2009learning .
-Given that most AT methods utilize WideResNetzagoruyko2016wide architectures, we select the commonly used WRN28-10 as the baseline network.
-During the entire training process, the learning rate is set to 0.01 and the number of sampling steps is 10.
-The initialization perturbation for adversarial samples is set to 8/255, and the number of adversarial sampling steps is 5.
-More details are shown in Appendix B.
-Note that EB-JDAT is a general optimization framework for JEM that is compatible with various JEM-based modelsgrathwohl2019your ; yang2021jem++ ; yang2023towards ; yang2023m .
-Therefore, we integrate EB-JDAT with SADAJEMyang2023towards and JEM++yang2021jem++ , two faster and more stable JEMs, to achieve improved performance.
+We conduct analysis and comparison experiments on several standard classifier datasets, including CIFAR-10 [20], CIFAR-100 [20] and ImageNet [8].
+Given that most AT methods utilize WideResNet [43] architectures, we select the commonly used WRN28-10 as the baseline network.
+All experiments were carried out on 3090 GPU.
+During the entire training process, we follow the default settings and SGLD hyperparameters of JEMs [39, 41], and the learning rate is set to 0.01.
+The initialization perturbation for adversarial samples is set to 8/255, and adversarial sampling steps is 5, while the perturbations are constrained using an $\ell_{\infty}$ norm.
+Note that EB-JDAT is a general optimization framework for JEM that is compatible with various JEM-based models [13, 39, 41, 40].
+Therefore, we integrate EB-JDAT with SADAJEM [41] and JEM++ [39], two faster and more stable JEMs, to achieve improved performance.
 4.2 Comparison Experiments And Analysis
 Compared with AT Methods
-We compare our approach with a range of AT methods, including five popular ones: PGD-AT mkadry2017towards , TRADES zhang2019theoretically , MART wang2019improving , AWP wu2020adversarial , and LBGAT cui2021learnable , as well as four SOTA methods: LAS-AWP jia2022adversarial , TDAT tong2024taxonomy , AGR-TRADES tong2024balancing , and VART zhao2024delving .
-All AT methods use the WRN34-10 network.
-For evaluation, we select the PGD mkadry2017towards and AutoAttack (AA) croce2020reliable .
+We compare our approach with a range of AT methods, including three popular ones: MART [34], AWP [37], and LBGAT [7], as well as six SOTA methods: LAS-AWP [19], UIAT [9], SGLR [23], TDAT [31], AGR-TRADES [30], and DHAT-CFA [45].
+For evaluation, we select the PGD-20 [25] and AutoAttack (AA) [6] which consists of APGD-CE [6], APGD-DLR [6], FAB [5] and Square [1].
+Both attacks utilizing the $\ell_{\infty}$ norm.
+All ATs are training on WRN28-10 [43].
 The results on CIFAR-10 and CIFAR-100 are shown in Tab.
-2, where EB-JDAT not only maintains the classification accuracy of clean data(92.28% on CIFAR-10; 70.19% on CIFAR-100) but also achieves robustness that even surpasses existing SOTA AT methods(+5.96% on CIFAR-10; +2.02% on CIFAR-100).
+2, where EB-JDAT not only maintains the classification accuracy of clean data (90.39% on CIFAR-10; 68.35% on CIFAR-100) but also achieves robustness that even surpasses existing SOTA AT methods (+6.38% on CIFAR-10; +1.23% on CIFAR-100).
+To evaluate the performance of our method on large-scale datasets, we further conducted comparative experiments on the ImageNet subset [8] in Tab.
+4.
+Given that adversarial training on the full ImageNet dataset requires substantial computational resources [38]), we selected a subset consisting of the first 100 categories from ImageNet-1000 and down-sampled the training images to a resolution of $64\times 64$.
 Thus, our method effectively balances the trade-off between accuracy on clean data and robustness against adversarial attacks.
-Furthermore, EB-JDAT is compatible with both SADAJEMyang2023towards and JEM++yang2021jem++ , thereby demonstrating the universality of our approach.
-Table 2: Camparative robustness (%) analysis with different AT methods.
+Furthermore, EB-JDAT is compatible with both SADAJEM [41] and JEM++ [39], thereby demonstrating the universality of our approach.
+Table 2: Comparative robustness (%) analysis with different AT methods. Results in bold indicate the best.
    CIFAR-10    CIFAR-100
    Method    Clean    PGD-20    AA    Clean    PGD-20    AA
-Standard Training 96.10 0 0 80.73 0 0
-PGD-AT 85.17 55.08 51.69 60.93 26.05 24.33
-TRADES 85.72 56.10 53.40 57.10 26.96 24.55
-MART 85.17 58.56 51.10 57.29 26.03 23.92
-AWP 85.57 58.13 53.90 60.38 33.86 28.86
-LBGAT 88.22 54.66 52.23 60.64 34.75 29.33
-LAS-AWP 87.74 60.16 55.52 64.89 36.36 30.77
-TDAT 82.25 56.03 54.06 57.32 33.17 26.61
-AGR-TRADES 86.50 52.56 52.56 58.17 28.34 25.84
-VART 84.21 58.32 54.24 / / /
-EB-JDAT-JEM++ 91.43 65.69 64.34 70.00 37.14 35.21
-EB-JDAT-SADAJEM 92.28 66.12 64.76 70.19 38.38 35.63
+   Standard Training    96.10    0.06    0    80.73    0    0
+   MART [34]    82.99    55.48    50.67    54.69    31.90    27.25
+   AWP [37]    82.67    57.21    51.90    57.94    33.75    28.90
+   LBGAT [7]    86.22    52.66    50.23    58.64    32.75    28.33
+   LAS-AWP [19]    87.74    60.16    55.52    64.89    36.36    30.77
+   UIAT [9]    82.94    58.12    52.17    57.65    33.91    29.03
+   SGLR [23]    85.72    56.10    53.40    61.02    32.98    28.50
+   TDAT [31]    82.25    56.03    54.06    57.32    33.17    26.61
+   AGR-TRADES [30]    86.50    52.56    52.56    58.17    28.34    25.84
+   DHAT-CFA [45]    84.49    62.38    54.05    61.54    37.15    30.93
+   EB-JDAT-JEM++    90.37    64.88    64.61    68.09    37.14    35.21
+   EB-JDAT-SADAJEM    90.39    68.76    66.30    68.35    38.38    35.63
+Table 3: Camparative robustness (%) analysis with AT methods that utilize generative models for data augmentation. Time is 3090 GPU hours. Results in bold indicate the best.
+  Data   Generate data   Epoch   Clean   AA   Time
+  CIFAR-10
+  SCORE [28]   1M   400   88.10   61.51   $\approx$ 1,438 h
+  The method proposed by [12]   100M   2000   87.50   63.38   $\approx$ 719,460 h
+  Better DM [35]   1M   400   91.12   63.35   $\approx$ 1,438 h
+  EB-JDAT-JEM++   n/a   100   90.37   64.61   31.66 h
+  EB-JDAT-SADAJEM   n/a   100   90.39   66.30   66.64 h
+  CIFAR-100
+  SCORE [28]   1M   400   62.08   31.40   $\approx$ 1,438 h
+  Better DM [35]   1M   400   68.06   35.65   $\approx$ 1,438 h
+  EB-JDAT-JEM++   n/a   100   68.09   35.21   31.66 h
+  EB-JDAT-SADAJEM   n/a   100   68.35   35.63   66.70 h
+Table 4: Comparison on ImageNet-Subset in terms of accuracy and robustness. Results in bold indicate the best.
+[…]
+MART [34] 49.30 21.84 17.40
+LAS-AT [19] 50.66 27.34 21.78
+WEAT [26] 61.90 30.27 24.52
+EB-JDAT-JEM++ 63.02 34.50 32.40
+Figure 3: Performance of generated samples on ImageNet subset (64x64) with EB-JDAT-JEM++.
+Table 5: Comparison of different hybrid models on CIFAR-10, IF (informative initialization). Results in bold indicate the best.
+[…]
+JEM variants
+JEM [13] 92.90 4.28 38.40 8.76
+JEM++ [39] 93.73 41.06 37.12 8.29
+SADAJEM [41] 96.03 29.63 17.38 8.07
+Energy-based AT
+JEAT [46] 85.16 28.43 38.24 8.80
+WEAT [26](PCA) 83.36 49.02 30.74 8.97
+WEAT [26](IF) 83.36 49.02 177.92 3.50
+EB-JDAT-JEM++ 90.37 64.61 39.67 7.66
+EB-JDAT-SADAJEM 90.39 66.30 27.42 8.05
+Figure 4: Performance comparison of generated samples from different methods, all methods are sampled by SGLD with informative initialization.
 Comparisons with AT Using Generated Data.
-Another strategy for improving robustness using generative models is to incorporate additional training data generated by generative models in AT pang2022robustness ; gowal2021improving .
-Therefore we compare EB-JDAT with SCORE pang2022robustness , Better DM wang2023better and the method proposed in gowal2021improving .
-SOCRE and Better DM utilize the SOTA Diffusion models ho2020denoising to generated addition training data, while gowal2021improving generate addition data from a ensemble of DDPM ho2020denoising , GANs goodfellow2014generative and VDVAE childvery generators.
-All methods are implemented using the WRN28-10 zagoruyko2016wide .
+Another strategy for improving robustness using generative models is to incorporate additional training data generated by generative models in AT [28, 12].
+Therefore we compare EB-JDAT with SCORE [28], Better DM [35] and the method proposed in [12].
+SOCRE and Better DM utilize the SOTA Diffusion models [15] to generated addition training data, while [12] generate addition data from a ensemble of DDPM [15], GANs [11] and VDVAE [3] generators.
+All methods are implemented using the WRN28-10 [43].
 The results are presented in Tab.
@@
 Although EB-JDAT does not rely on additional generated data or increased training epochs, it still achieves competitive performance in terms of clean accuracy and robustness.
-EB-JDAT even outperforms SCORE-1M, Better DM-1M and gowal2021improving .
-Unlike discriminative learning approaches pang2022robustness ; gowal2021improving ; wang2023better , which struggles to capture the underlying structure of both clean and adversarial data distributions, EB-JDAT, as a hybrid generative model, provides more accurate gradient estimations of clean/adversarial data densities, leading to substantial improvements in both accuracy and robustness.
+EB-JDAT even outperforms SCORE-1M and [12].
+Unlike discriminative learning approaches [28, 12, 35], which struggles to capture the underlying structure of both clean and adversarial data distributions, EB-JDAT, as a hybrid generative model, provides more accurate gradient estimations of clean/adversarial data densities, leading to substantial improvements in both accuracy and robustness.
 Notably, this is achieved without requiring additional training data and with minimal computational overhead.
-Table 3: Camparative robustness (%) analysis with AT methods that utilize generative models for data augmentation, P denotes EB-JDAT with perceptual loss.
-      Data       Generate data       Epoch       Clean       AA
-      CIFAR-10
-      SCORE pang2022robustness       1M       400       88.10       61.51
-      The method proposed by gowal2021improving       100M       2000       87.50       63.38
-      Better DM wang2023better       1M       400       91.12       63.35
-      Better DM wang2023better       50M       1600       92.27       67.17
-      EB-JDAT-SADAJEM(P)       n/a       100       92.28       64.76
-      EB-JDAT-SADAJEM       n/a       100       89.79       66.30
-      CIFAR-100
-      SCORE pang2022robustness       1M       400       62.08       31.40
-      Better DM wang2023better       1M       400       68.06       35.65
-      Better DM wang2023better       50M       1600       72.58       38.83
-      EB-JDAT-SADAJEM(P)       n/a       100       70.19       35.63
-      EB-JDAT-SADAJEM       n/a       100       68.35       38.46
 Compared with JEMs and Energy-based AT.
-In this section, we compare EB-JDAT with other JEMs grathwohl2019your ; yang2021jem++ ; yang2023towards and energy-based AT zhu2021towards ; mirza2024shedding in terms of classification accuracy, robustness and generative performance.
-Sec.
-4.2 presents an evaluation of our method against JEM variants with respect to classification accuracy and generation quality.
+In this section, we compare EB-JDAT with other JEMs [13, 39, 41] and energy-based AT [46, 26] in terms of classification accuracy, robustness and generative performance.
+Tab.
+5 presents an evaluation of our method against JEM variants with respect to classification accuracy and generation quality.
 To assess generation quality, we employ the Fréchet Inception Distance (FID) and Inception Score (IS) as metrics.
 Compared to JEMs, EB-JDAT maintains comparable accuracy and generative performance, while EB-JDAT-SADAJEM even outperforms JEM (FID-10.98) and JEM++ (FID-9.7) in terms of generation.
-Moreover, EB-JDAT exhibits significantly improved robustness over JEMs, as shown in Fig.
+Moreover, EB-JDAT exhibits significantly improved robustness over JEMs, although it exhibits a slight decrease in classification accuracy and generative ability relative to the strongest JEMs [40], as shown in Tab.
+5.
+This reflects an intrinsic trade-off: adversarial training uses adversarial examples to regularize decision boundaries and reshape the energy landscape, which can shift probability mass away from the clean data manifold and slightly reduce sample fidelity, which are shown in Fig.
 2.
-Thus, EB-JDAT bridges the gap between classification, robustness, and generation.
-Table 4: Performance comparison of different hybrid models on CIFAR-10, IF denotes informative initialization.
-[…]
-JEM variants
-JEM grathwohl2019your 92.90 38.40 8.76
-JEM++ yang2021jem++ 93.73 37.12 8.29
-SADAJEM yang2023towards 96.03 11.40 8.63
-Energy-based AT
-JEAT zhu2021towards 85.16 38.24 8.80
-WEAT mirza2024shedding (PCA) 83.36 30.74 8.97
-WEAT mirza2024shedding (IF) 83.36 177.92 3.50
-EB-JDAT-JEM++ 91.43 38.89 7.66
-EB-JDAT-SADAJEM 92.28 27.42 8.05
-Figure 2: Comparative robustness (%) analysis with different energy-based AT.
-Next, we compare EB-JDAT with energy-based AT zhu2021towards ; mirza2024shedding .
+Overall, compared with existing JEMs [13, 39, 41], EB-JDAT achieves the best balance among classification, generation, and robustness.
+Next, we compare EB-JDAT with energy-based AT [46, 26].
 First, we compare the accuracy and robustness with them, which is reported in Fig.
-2.
+1.
 The results demonstrate that our method significantly enhances robustness with minimal accuracy degradation, outperforming existing energy-based AT.
-Moreover, our method yields superior generative performance compared to energy-based methods, which are shown in Sec.
-4.2.
+Moreover, our method yields superior generative performance compared to energy‐based methods, which are shown in Tab.
+5.
 EB-JDAT surpasses JEAT and WEAT with an FID reduction of 10.82 and 3.32, respectively.
-Unlike JEAT zhu2021towards directly introduces adversarial samples into the JEMs training process to model the joint distribution ${p}_{\theta}(\tilde{\mathbf{x}},y)$, we model the joint probability $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$ to capture the distribution of clean and adversarial samples more accurately, thus performing better in accuracy, robustness and generative ability.
-On top of that, WEATmirza2024shedding reinterprets TRADESzhang2019theoretically as an EBM to model the conditional distribution $p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$, which is still a discriminative model.
-Although mirza2024shedding notes that PCA-based initialization can give WEAT generative capabilities(FID=30.74), but under the same comparison of informative initialization, WEAT shows poor generative performance(FID=177.92).
+Unlike JEAT [46] directly introduces adversarial samples into the JEMs training process to model the joint distribution ${p}_{\theta}(\tilde{\mathbf{x}},y)$, we model the joint probability $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$ to capture the distribution of clean and adversarial samples more accurately, thus performing better in accuracy, robustness and generative ability.
+On top of that, WEAT [26] reinterprets TRADES [44] as an EBM to model the conditional distribution $p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$, which is still a discriminative model.
+Although [26] notes that PCA-based initialization can give WEAT generative capabilities (FID=30.74), but under the same comparison of informative initialization, WEAT shows poor generative performance (FID=177.92).
 Consequently, EB-JDAT outperforms previous energy-based AT in classification accuracy, robustness, and generative capability.
@@
 To further visually explore generative ability of EB-JDAT, we generate images via SGLD sampling with informative initialization for conventional AT-trained models, energy-based AT-models, JEMs, and EB-JDAT in Fig.
-3.
+4.
 As demonstrated in Fig.
-3, the conventional AT-trained models(a, b) and energy-based AT-models(e) possess limited generative capacity, resulting in difficulties in generating high-quality images.
+4, the conventional AT-trained models (a, b) and energy-based AT-models (e) possess limited generative capacity, resulting in difficulties in generating high-quality images.
 Conversely, EB-JDAT produce higher-resolution images with richer background details, which is more closely matching the features of the original dataset.
 In addition, EB-JDAT (g, h) surpasses JEM (f) and perform competitively with JEM++ and SADAJEM in generative capacity, producing more diverse and detailed images with only a slight reduction in sharpness, particularly in the ’car’ category.
-Figure 3: Performance comparison of generated samples from different methods, all methods are sampled by SGLD with informative initialization.
+We also implement our method on the ImageNet subset and generate higher resolution images, as shown in Fig.
+3.
 4.3 Ablation Study
-Table 5: Ablation study of EB-JDAT-JEM++, ECO denotes the epoch at which collapse occurs.
+Ablation of Weighted Gradient Components.
+In EB-JDAT, we employ a weighted combination of gradients, as detailed in Algorithm 1, to compute the overall optimization gradient, expressed as $h_{\theta}=w_{1}h_{1}+w_{2}h_{2}+w_{3}h_{3}$.
+To assess the individual contribution of each weight, we perform ablation studies by setting each of $w_{1}$, $w_{2}$, and $w_{3}$ to zero and evaluating their impact on overall performance, including (1) standard AT (NO.1) , (2) jointly modeling the energy distributions of adversarial and clean samples (NO.2) , and (3) jointly modeling the energy distributions of adversarial, clean, and generated samples (NO.3 and NO.4).
+The results, presented in Table 6, reveal two key findings: (1) $\frac{\partial\log p_{\theta}(\tilde{{\mathbf{x}}}|{\mathbf{x}})}{\partial\theta}$ ($h_{2}$) is not only the key to bridge the gap of accuracy, generation and robustness, but also stabilizes the training phase, reducing the risk of model collapse.
+(2) $\frac{\partial\log p_{\theta}({\mathbf{x}})}{\partial\theta}$ ($h_{1}$) can improve the ability of model in both classification and generation.
+Consequently, experimental results indicate that $w_{1}=1$, $w_{2}=1$, and $w_{3}=1$ provides the best trade-off between accuracy, robustness, generative capability, and stability.
+This configuration is therefore adopted as the default setting.
+Additionally, reducing $w_{2}$ can improve classification accuracy and generative quality, albeit with a slight decrease in robustness.
+Ablation of Adversarial Sampling Steps.
+Further, we ablate the number of adversarial sampling steps K $\in${1,5,10,20} while fixing the perturbation budget and step size.
+As shown in Fig.
+5, a setting of 5 steps provided the best trade-off, allowing the model to converge stably.
+This choice was made to balance two factors: finding sufficiently challenging adversarial examples and avoiding model collapse, which is a known issue for energy-based models on complex data, as we note in our limitations section .
+Table 6: Ablation study of EB-JDAT-JEM++, ECO denotes the epoch at which collapse occurs. Results in bold indicate the best.
 […]
-1 0 0.1 0.9 90.61 66.41 64.71 146.02 4.44 n/a
-2 0.5 0 0.5 74.16 46.28 43.56 69.37 5.95 20
-3 0.9 0.1 0 12.59 11.59 11.48 74.29 5.17 44
-4 0.35 0.25 0.35 87.13 60.31 58.20 48.52 6.56 70
-5 0.45 0.05 0.45 91.43 65.69 64.34 38.89 7.66 n/a
-In EB-JDAT, we employ a weighted combination of gradients, as detailed in Algorithm 1, to compute the overall optimization gradient, expressed as $h_{\theta}=w_{1}h_{1}+w_{2}h_{2}+w_{3}h_{3}$.
-To assess the individual contribution of each weight, we perform ablation studies by setting each of $w_{1}$, $w_{2}$, and $w_{3}$ to zero and evaluating their impact on overall performance.
-The results, presented in Table 5, reveal two key findings: (1) When $w_{2}=0$, the optimization focuses solely on classification loss and the clean-sample distribution.
-Although this setting enhances robustness, the significant energy gap between clean and adversarial samples prevents model convergence, leading to model collapse by the 20th epoch.
-(2) Maintaining a small value for $w_{2}$ not only balances accuracy and robustness but also stabilizes the training phase, reducing the risk of model collapse.
-Consequently, experimental results indicate that $w_{1}=0.45$, $w_{2}=0.05$, and $w_{3}=0.45$ yields optimal performance in terms of accuracy, robustness, generative capability, and stability.
-This configuration is therefore adopted as the default setting.
-5 Limitation And Broader Impact
+1 0 0 1 88.95 62.96 173.53 41
+2 0 1 1 89.84 64.69 42.57 n/a
+3 1 0.5 1 90.39 64.09 40.12 n/a
+4 1 1 1 90.37 64.61 39.67 n/a
+Figure 5: Ablation study of adversarial sampling steps. ECO=100 indicates no collapse. Time denotes the per-epoch cost.
+5 Limitation And Discussion
 Training EB-JDAT on complex, high-dimensional data remains challenging.
-This challenge is also encountered by JEMs,including JEM grathwohl2019your , JEM++ yang2021jem++ and SADAJEM yang2023towards .
+This challenge is also encountered by JEMs,including JEM [13], JEM++ [39] and SADAJEM [41].
 This instability arises from the typically sharp probability distribution of real data in high-dimensional space, which leads to inaccurate guidance for image sampling in regions with low data density.
 Although the approach we propose is a general and flexible optimization framework for all JEMs, considering training stability, we recommend training within faster and more stable JEM variants, such as JEM++ and SADAJEM.
-Nevertheless, our approach allows a standard classifier to achieve good generative ability and robustness without sacrificing much accuracy, resulting in the best overall performance in terms of accuracy, generative capability, and robustness.
-From a societal perspective, a robustness generative-classifier more effectively mitigates data scarcity and adversarial perturbations.
-However, if the training dataset itself is high-risk, the model may generate equally high-risk content, potentially posing harm to society.
-To minimize such harm, we will soon release user guidelines and begin implementing safety filters.
+Nevertheless, our method significantly enhances the robustness of JEMs, surpassing SOTA AT [45], while incurring only a slight degradation in accuracy and generative performance, thereby achieving the best overall trade-off among robustness (68.76%) , accuracy (90.39%) , and generation (FID=27.42) .
 6 Conclusion
-In this work, we systematically investigate the underlying causes of the performance gap between discriminate classifiers, robust classifiersmkadry2017towards ; zhang2019theoretically and JEM-based classifiers grathwohl2019your ; yang2023towards in terms of accuracy, generative capability, and adversarial robustness, and propose a general and flexible optimization framework, called EB-JDAT to eliminate the inherent trilemma.
+In this work, we systematically investigate the underlying causes of the performance gap between discriminate classifiers, robust classifiers [25, 44] and JEM-based classifiers [13, 41] in terms of accuracy, generative capability, and adversarial robustness, and propose a general and flexible optimization framework, called EB-JDAT to eliminate the inherent trilemma.
 Our experiments validate the effectiveness of these techniques across various benchmarks, demonstrating superior performance in achieving the triple trade-off, with particularly impressive SOTA results in robustness.
-Appendix A Ethics Statement
-We exclusively employed publicly available datasets and models for research purposes and did not involve human subjects.
-EB-JDAT is a robust generative discriminator with powerful image-generation capabilities; if misused, it could produce content that poses societal risks, and no safeguards currently limit its outputs.
-Nevertheless, our approach achieves significantly greater robustness than earlier generative discriminators, ensuring its safe deployment in real-world applications.
-We therefore contend that the benefits of our work far outweigh any potential harms arising from disclosure of our methods.
-In addition, a fundamental classifier endowed with generative capacity and discriminative robustness more effectively mitigates data scarcity and adversarial perturbations.
-Appendix B Experimental details
-The experiments were carried out on a platform featuring an AMD EPYC 7542 32-core CPU with a clock speed of 2039.813 GHz and an NVIDIA GeForce RTX 3090 GPU, equipped with 24 GB of RAM.
-The proposed method was implemented using the open-source PyTorch machine learning framework.
-In addition, we use the same hyperparameter settings and validation-set configuration for training and evaluation as in JEM variants [2, 3].